OpenStudy (anonymous):

Simplify. Please be sure to show all of your work. 12-1/9÷4/3+(1/2)^2

OpenStudy (anonymous):

73/6

OpenStudy (anonymous):

= 12 - 1/12 + 1/4 = 73/6

OpenStudy (trancenova):

lots of working (now where did I put that calculator!): 12 - (1/9)/(4/3) + (1/2)^2 = 12 - (1/9)*(3/4) + (1^2/2^2) = 12 - (1*3)/(9*4) + (1/4) = 12 - 3/36 + 1/4 (lowest common demoninator of 4 and 36) = 12 - 3/36 + (9*1)/(4*9) = 12 - 3/36 + 9/36 = 12 + 6/36 = 12 + 1/6 = 73/6

jhonyy9 (jhonyy9):

check the denominator common for 1 4 1 (12)(36) - 4 + 4(12) + 9 432 -4 +48 +9 485 12 - --- + --- + --- = ---------------------- = --------------- = ----- 9 3 4 36 36 36

jhonyy9 (jhonyy9):

than after 1/9 you have wrote divide sign - if this is true like , hence - you will get 1 3 1 1 1 1 + 3 1 35 12 - --- * ---- + ---- = 12 - ---- + --- = 12 - ------ = 12 - ---- = ---- 9 4 2^2 12 4 12 3 3